# Deep Active Speech Cancellation with Multi-Band Mamba Network

### v1

@@ Abstract
We present a novel deep learning network for Active Speech Cancellation (ASC), advancing beyond Active Noise Cancellation (ANC) methods by effectively canceling both noise and speech signals. The proposed Multi-Band Mamba architecture segments input audio into distinct frequency bands, enabling precise anti-signal generation and improved phase alignment across frequencies. Additionally, we introduce an optimization-driven loss function that provides near-optimal supervisory signals for anti-signal generation. Experimental results demonstrate substantial performance gains, achieving up to 7.2dB improvement in ANC scenarios and 6.2dB in ASC, significantly outperforming existing methods. Audio samples are available at this link.
@@ 1 Introduction
Active Noise Cancellation (ANC) is a critical audio processing technique aimed at eliminating unwanted noise by generating an anti-noise signal (Lueg, 1936; Nelson & Elliott, 1991; Fuller et al., 1996; Hansen et al., 1997; Kuo & Morgan, 1999). ANC has practical applications in improving hearing devices for individuals with hearing impairments and reducing chronic noise exposure, thereby mitigating hearing loss risks. It also enhances focus, productivity, and listening experiences while reducing stress. Traditional ANC algorithms, like LMS and its deep learning variants (Zhang & Wang, 2021; Park et al., 2023; Mostafavi & Cha, 2023; Cha et al., 2023; Pike & Cheer, 2023; Singh et al., 2024), have been widely adopted. However, these methods face limitations when dealing with more complex and high-frequency audio signals, as they are primarily designed to target noise. This paper addresses Active Speech Cancellation (ASC), which expands upon ANC by targeting the cancellation of both noise and speech signals. To our knowledge, this is the first work to actively cancel both noise and speech using deep learning, setting it apart from existing methods and enabling new research directions.
We propose a novel Multi-Band Mamba architecture, which partitions input signals into frequency bands, enabling precise anti-signal generation for improved phase alignment and cancellation. This design is particularly effective for speech signals, as it accounts for their broader frequency spectrum, including high frequencies often under-addressed by other methods. Coupled with an optimization-driven loss function, this approach achieves improved performance in dynamic acoustic scenarios. Results demonstrate up to a 7.2 dB improvement in ANC and a 6.2 dB gain in ASC for speech signals, outperforming deep-learning based baselines, which are considered state-of-the-art in the field.
@@ 2 Related Work
Active Noise Cancellation: The concept of ANC was first introduced by Lueg (Lueg, 1936), focusing on sound oscillation cancellation. Given that ANC algorithms must adapt to variations in amplitude, phase, and noise source movement (Nelson & Elliott, 1991; Fuller et al., 1996; Hansen et al., 1997; Kuo & Morgan, 1999), most ANC algorithms are based on the Least Mean Squares (LMS) algorithm (Burgess, 1981), which is effective in echo cancellation. The FxLMS (Filtered-x LMS) algorithm extends LMS by using an adaptive filter to account for distortions in the primary and secondary paths. Boucher et al. (1991) analyzed errors in FxLMS due to inaccuracies in estimating the secondary path inverse, where nonlinear distortions degrade performance. Several methods address these issues: the Filtered-S LMS (FSLMS) (Das & Panda, 2004) algorithm uses a Functional Link Artificial Neural Network (FLANN) (Patra et al., 1999) to handle nonlinearity, while the Volterra Filtered-x LMS (VFXLMS) (Tan & Jiang, 2001) employs a multichannel structure. The Bilinear FxLMS (Kuo & Wu, 2005) improves nonlinearity modeling, and the Leaky FxLMS (Tobias & Seara, 2005) introduces a leakage term to mitigate overfitting. The Tangential Hyperbolic Function-based FxLMS (THF-FxLMS) (Ghasemi et al., 2016) models saturation effects for enhanced performance. Gannot & Yeredor (2003) proposed blind source separation for noise cancellation. Moreover, Oppenheim et al. (1994) proposed single channel ANC based on Kalman filter formulation (Revach et al., 2021) and Rafaely (2009) investigated spherical loudspeaker arrays for local sound control.
@@
ANC using deep learning was first proposed by Zhang & Wang (2021) with a convolutional-LSTM network for estimating both amplitude and phase of the canceling signal $y(n)$. Recurrent CNNs were later explored by Park et al. (2023); Mostafavi & Cha (2023); Cha et al. (2023), and autoencoder-based networks (Singh et al., 2024), along with fully connected neural networks, were also applied to the problem (Pike & Cheer, 2023). Shi et al. (2020, 2022b); Luo et al. (2022); Park & Park (2023); Shi et al. (2023a); Luo et al. (2023b, a, 2024b) have developed methods that select fixed-filter ANC (SFANC) from pre-trained control filters to achieve fast response times. Concurrently, Zhu et al. (2021); Shi et al. (2022a); Zhang & Wang (2023); Shi et al. (2023b); Antoñanzas et al. (2023); Xiao et al. (2023); Zhang et al. (2023b); Shi et al. (2024) advanced multichannel ANC systems. Luo et al. (2023c) introduced a CNN-based approach for real-time ANC, further enhanced with Kalman filtering. Zhang et al. (2023a) incorporated an attention mechanism for real-time ANC using the Attentive Recurrent Network (ARN)(Pandey & Wang, 2022). Other significant real-time ANC contributions include genetic and bee colony algorithm-based methods (Ren & Zhang, 2022; Zhou et al., 2023).
@@
Active Speech Cancellation: ASC has been explored in various studies, each employing different approaches to predict and cancel unwanted speech signals. Kondo & Nakagawa (2007) introduced an ASC method using a Linear Predictive Coding (LPC) model to predict the speech signal for generating the canceling signal $y(n)$. Donley et al. (2017) took a different approach by controlling the sound field to cancel speech using a linear dipole array of loudspeakers and a single microphone, effectively reducing the speech signal in the target area. Iotov et al. (2022) employed a long-term linear prediction filter to anticipate incoming speech, enabling the cancellation of the speech signal. Additionally, Iotov et al. (2023) proposed HOSpLP-ANC, which combines a high-order sparse linear predictor with the LMS algorithm for effective speech cancellation.
@@
Mamba architecture: Recently, the Mamba architecture has been introduced (Gu & Dao, 2023; Dao & Gu, 2024), leveraging State Space Models (SSMs) to achieve notable improvements in various audio-related tasks. One of the key advantages of the Mamba architecture is its ability to perform fast inference, especially when handling sequences up to a million in length, which represents a significant improvement over traditional generative architectures. This has enabled advancements in several applications, including automatic speech recognition (Zhang et al., 2024b, a), speech separation (Jiang et al., 2024a; Li & Chen, 2024), speech enhancement (Chao et al., 2024; Luo et al., 2024a; Quan & Li, 2024), speech super-resolution (Lee & Kim, 2024), sound generation (Jiang et al., 2024b), audio representation (Shams et al., 2024; Yadav & Tan, 2024; Erol et al., 2024), sound localization (Xiao & Das, 2024; Mu et al., 2024), audio tagging (Lin & Hu, 2024), and deepfake audio detection (Chen et al., 2024).
Figure 1: Typical ANC system diagram.
3 Method
Backgorund
@@ 3 Method
The signal processing framework of a typical feedforward ANC system is detailed, emphasizing the roles of the primary and secondary acoustic paths. In such systems, reference and error microphone are utilized to generate a canceling signal that minimizes unwanted noise. The primary path $P(z)$ represents the acoustic transfer function from the noise source to the error microphone, while the secondary path $S(z)$ represents the acoustic transfer function from the loudspeaker to the error microphone. The signal captured by the reference microphone is denoted as $x(n)$, while the signal captured by the error microphone is denoted as $e(n)$. These signals are fed into the ANC controller, which generates a canceling signal $y(n)$. The canceling signal is then played through a loudspeaker, referred to as $f_{LS}$, producing $f_{LS}\{y(n)\}$, which aims to suppress the unwanted noise near the error microphone. The loudspeaker output $f_{LS}\{y(n)\}$, after passing through the secondary path $S(z)$, generates the anti-signal denoted by $a(n)$. The relationship is described by the following equation:
$a(n)=S(z)\ast f_{LS}\{y(n)\}$ (1)
Similarly, the reference signal $x(n)$, transmitted through the primary path $P(z)$, produces the primary signal denoted by $d(n)$, which is expressed as:
$d(n)=P(z)\ast x(n)$ (2)
The error signal $e(n)$, defined as the difference between the primary signal $d(n)$ and the anti-signal $a(n)$, is defined as:
@@
where u and v are the vector representations of the signals $u(n)$ and $v(n)$ such that $\textbf{u}=[u(1),...,u(M)]$ and $\textbf{v}=[v(1),...,v(M)]$. Here, $M$ represents the total number of samples. Typically, $u(n)$ refers to the target signal, while $v(n)$ denotes the estimated signal. A lower NMSE value indicates a better estimation, reflecting a closer alignment between the estimated signal and the target signal. In the context of ANC, typically $u(n)$ is the primary signal $d(n)$, while $v(n)$ will be the anti-signal $a(n)$. A schematic representation of the ANC system is illustrated in Figure 1. Appendix A provides an explanation of the acoustic dynamics of ANC systems in in-ear headphones.
Method
The proposed method utilizes a novel architecture that integrates the Mamba framework (Gu & Dao, 2023) for the generation of the anti-signal. The architecture features a filter bank that decomposes the input signal into multiple frequency bands, each processed by an encoder and a masking Mamba network, incorporating Dual-path Mamba blocks (Jiang et al., 2024a) for enhanced feature extraction. The outputs of the multi-band masking are then concatenated and passed through a decoder. Furthermore, we introduce a new loss function that leverages the near-optimal anti-signal as the ground truth, significantly improving the precision of the anti-signal generation process. A diagram of the proposed architecture is shown in Fig.2.
Figure 2: DeepASC Architecture: The reference signal is decomposed into frequency bands, encoded, processed by Mamba-Band blocks, concatenated, and decoded to reconstruct the signal.
@@ 3.1 DeepASC Architecture
Let $x(n)$ be the reference signal such that $1\leq n\leq M$. The reference signal $x(n)$ is decomposed into $Q\in\mathbb{N}$ different frequency bands $x_{1}(n),\ldots,x_{Q}(n)$. These frequency bands are evenly divided such that for the maximum frequency $F$, the $i$-th frequency band $x_{i}(n)$ covers the frequency range $\left[(i-1)\frac{F}{Q},i\frac{F}{Q}\right]$ where $1\leq i\leq Q$. In addition to the decomposed bands, the original full-band signal $x(n)$ is included as $x_{0}(n)$. Each band $x_{i}(n)$ (where $0\leq i\leq Q$, the zero index is for the entire unfiltered band) is then processed through its own Mamba-Band block (MB-block). Each MB-block comprises an encoder and a masking network that utilize Mamba-based layers. Within each MB-block, the encoder consists of a one-dimensional convolution layer $E_{i}$ with a kernel size $k$ and a stride of $k/2$. The encoder transforms the $i$-th reference signal $x_{i}(n)$ into a two-dimensional latent representation:
@@ 3.2 Optimization Objective
The training protocol for the proposed method consists of two distinct phases: (i) ANC loss minimization, and (ii) near optimal anti-signal optimization. Each phase employs the NMSE loss function (Eq. 4) but with different optimization objectives.
@@
Near Optimal Anti-Signal Optimization (NOAS): A key challenge in formulating ANC as a supervised learning problem is defining a training objective that accounts for the characteristics of the secondary path $S(z)$ and the primary path $P(z)$. In an ANC algorithms, the output $y(n)$ is processed by a nonlinearity function $f_{LS}$ and then propagated through the secondary path $S(z)$. The training objective aims to minimize the error signal $e(n)$, which represents the residual noise after cancellation.
However, this process becomes problematic when the secondary path $S(z)$ attenuates certain frequencies that are present in the primary signal $d(n)$. Under the vanilla loss function (e.g., Eq. 10), the model can be unfairly penalized for high error signals in these attenuated frequency bands, even when it has generated an optimal anti-signal. This occurs because the secondary path inherently suppresses these frequencies, leading to residual energy in the error signal $e(n)$. As a result, the training process encounters discrepancies that hinder the model’s ability to learn effectively.
To address this challenge, we propose the NOAS loss function. The NOAS loss symmetrically incorporates the secondary path $S(z)$ on both sides of the NMSE calculation. By doing so, it ensures that any frequencies nullified by $S(z)$ are also excluded from the target, thereby mitigating the contribution of these frequencies to the error signal. Specifically, each reference signal $x(n)$ is associated with its NOAS target $y^{*}(n)$. To determine the near-optimal anti-signal $y^{*}(n)$, we employ a gradient descent-based algorithm during a pre-processing stage. This stage operates over each example, solving the following optimization problem for each reference signal $x(n)$:
@@
Note that the optimization occurs in the S-projected space, rather than directly in the canceling signal space (i.e. $\text{NMSE}\left[\textbf{y}^{*},\textbf{y}\right]$). The S projection ensures consistency and leverages prior knowledge from the initial training phase. Furthermore, it is necessitated by the nature of convolution, which is not a injective operation. As a result, multiple distinct instances, denoted as $\mathbf{y}_{1},\dots,\mathbf{y}_{n}$, may satisfy $\mathbf{S}\ast f_{LS}\{\mathbf{y}_{1}\}=\dots=\mathbf{S}\ast f_{LS}\{\mathbf{y}_{n}\}$. Consequently, even if a trained model satisfies $\mathbf{S}\ast f_{LS}\{\mathbf{y}\}\approx\mathbf{P}\ast\mathbf{x}$, it does not necessarily follow that $\mathbf{y}\approx\mathbf{y}^{*}$. This phenomenon is further corroborated by the results in Table 1, which report NMSE measurements on audio samples from the train set. Prior to NOAS fine-tuning, the NMSE between $\left[\mathbf{y}^{*},\mathbf{y}\right]$ is $-9.85$ dB, while the NMSE between $\left[\mathbf{P}\ast\mathbf{x},\mathbf{S}\ast\mathbf{y}\right]$ and $\left[\mathbf{S}\ast\mathbf{y}^{*},\mathbf{S}\ast\mathbf{y}\right]$ are $-16.53$ dB and $-18.53$ dB, respectively.
Notably, the NMSE between $\left[\mathbf{S}\ast\mathbf{y}^{*},\mathbf{S}\ast\mathbf{y}\right]$ is the lowest. This result aligns with the intuition that $\mathbf{S}\ast\mathbf{y}^{*}$ serves as feasible optimization target for $\mathbf{S}\ast\mathbf{y}$, unlike $\mathbf{P}\ast\mathbf{x}$, which is constrained by frequency limitations, as discussed earlier. A visual intuition for this phenomenon is provided in Figure 3.
Table 1: Comparison of NMSE distances for different objectives, with and without NOAS optimization.
@@
- NOAS -9.85 -16.53 -18.56
+ NOAS -12.77 -17.60 -19.62
Figure 3: Schematic representation of signal transformations via the secondary path.
(a) Engine
(b) Babble
(c) Factory
(d) Speech
Figure 4: Comparison of NMSE over time for different noise types.
Table 2: Average NMSE for DeepASC and other algorithms across various noise types and nonlinear distortions.

### v2

# Deep Active Speech Cancellation with Mamba-Masking Network

@@ Abstract
We present a novel deep learning network for Active Speech Cancellation (ASC), advancing beyond Active Noise Cancellation (ANC) methods by effectively canceling both noise and speech signals. The proposed Mamba-Masking architecture introduces a masking mechanism that directly interacts with the encoded reference signal, enabling adaptive and precisely aligned anti-signal generation—even under rapidly changing, high-frequency conditions, as commonly found in speech. Complementing this, a multi-band segmentation strategy further improves phase alignment across frequency bands. Additionally, we introduce an optimization-driven loss function that provides near-optimal supervisory signals for anti-signal generation. Experimental results demonstrate substantial performance gains, achieving up to 7.2dB improvement in ANC scenarios and 6.2dB in ASC, significantly outperforming existing methods.
@@ 1 Introduction
Active Noise Cancellation (ANC) is a critical audio processing technique aimed at eliminating unwanted noise by generating an anti-noise signal [1, 2, 3, 4, 5]. ANC has practical applications in improving hearing devices for individuals with hearing impairments and reducing chronic noise exposure, thereby mitigating hearing loss risks. It also enhances focus, productivity, and listening experiences while reducing stress. Traditional ANC algorithms, like LMS and its deep learning variants [6, 7, 8, 9, 10, 11], have been widely adopted. However, these methods face limitations when dealing with more complex and high-frequency audio signals, as they are primarily designed to target noise. This paper addresses Active Speech Cancellation (ASC), which expands upon ANC by targeting the cancellation of both noise and speech signals. To our knowledge, this is the first work to actively cancel both noise and speech using deep learning, setting it apart from existing methods and enabling new research directions.
We propose a novel Mamba-Masking multi-band architecture that applies a masking mechanism to the encoded signal. This facilitates precise anti-signal generation, enhancing phase alignment and improving ANC performance. This design is particularly effective for speech signals, as it accounts for their broader frequency spectrum. Coupled with an optimization-driven loss function, this approach achieves improved performance in dynamic acoustic scenarios. Results demonstrate up to a 7.2 dB improvement in ANC and a 6.2 dB gain in ASC for speech signals, outperforming deep-learning based baselines, which are considered state-of-the-art in the field.
@@ 2.1 Active Noise Cancellation
The concept of ANC was first introduced by Lueg [1], focusing on sound oscillation cancellation. Given that ANC algorithms must adapt to variations in amplitude, phase, and noise source movement [2, 3, 4, 5], most ANC algorithms are based on the Least Mean Squares (LMS) algorithm [12], which is effective in echo cancellation. The FxLMS algorithm extends LMS by using an adaptive filter to correct distortions in primary and secondary paths. Boucher et al. [13] examined errors in FxLMS due to inaccuracies in estimating the secondary path inverse, where nonlinearities affect performance. Solutions such as the Filtered-S LMS (FSLMS) [14], which uses a Functional Link Artificial Neural Network (FLANN) [15], and the Volterra Filtered-x LMS (VFXLMS) [16], which employs a multichannel structure, address these issues. The Bilinear FxLMS [17] improves nonlinearity modeling, and the Leaky FxLMS [18] introduces a leakage term to mitigate overfitting. The Tangential Hyperbolic Function-based FxLMS (THF-FxLMS) [19] models saturation effects for enhanced performance. Gannot and Yeredor [20] proposed blind source separation for noise cancellation. Moreover, Oppenheim et al. [21] proposed single channel ANC based on Kalman filter formulation [22] and Rafaely [23] investigated spherical loudspeaker arrays for local sound control.
@@
ANC using deep learning was first proposed by Zhang and Wang [6] with a convolutional-LSTM network for estimating both amplitude and phase of the canceling signal $y(n)$. Recurrent CNNs were later explored by Park et al. [7], Mostafavi and Cha [8], Cha et al. [9], and autoencoder-based networks [11], along with fully connected neural networks, were also applied to the problem [10]. Shi et al. [24, 25], Luo et al. [26], Park and Park [27], Shi et al. [28], Luo et al. [29, 30, 31] have developed methods that select fixed-filter ANC (SFANC) from pre-trained control filters to achieve fast response times. Concurrently, Zhu et al. [32], Shi et al. [33], Zhang and Wang [34], Shi et al. [35], Antoñanzas et al. [36], Xiao et al. [37], Zhang et al. [38], Shi et al. [39] advanced multichannel ANC systems. Luo et al. [40] introduced a CNN-based approach for real-time ANC, further enhanced with Kalman filtering. Zhang et al. [41] incorporated an attention mechanism for real-time ANC using the Attentive Recurrent Network (ARN)[42]. Other significant real-time ANC contributions include genetic and bee colony algorithm-based methods [43, 44].
2.2 Active Speech Cancellation
@@
ASC has been explored in various studies, each employing different approaches to predict and cancel unwanted speech signals. Kondo and Nakagawa [45] introduced an ASC method using a Linear Predictive Coding (LPC) model to predict the speech signal for generating the canceling signal $y(n)$. Donley et al. [46] took a different approach by controlling the sound field to cancel speech using a linear dipole array of loudspeakers and a single microphone, effectively reducing the speech signal in the target area. Iotov et al. [47] employed a long-term linear prediction filter to anticipate incoming speech, enabling the cancellation of the speech signal. Additionally, Iotov et al. [48] proposed HOSpLP-ANC, which combines a high-order sparse linear predictor with the LMS algorithm for effective speech cancellation.
2.3 Mamba Architecture
@@
Recently, the Mamba architecture has been introduced [49, 50], leveraging State Space Models (SSMs) to achieve notable improvements in various audio-related tasks. One of the key advantages of the Mamba architecture is its ability to perform fast inference, especially when handling sequences up to a million in length, which represents a significant improvement over traditional generative architectures. This has enabled advancements in several applications, including automatic speech recognition [51, 52], speech separation [53, 54], speech enhancement [55, 56, 57], speech super-resolution [58], sound generation [59], audio representation [60, 61, 62], sound localization [63, 64], audio tagging [65], and deepfake audio detection [66].
3 Background
@@
The feedforward ANC system consists of reference and error microphones, a loudspeaker, and two acoustic transfer paths: the primary path $P(z)$, from the noise source to the error microphone, and the secondary path $S(z)$, from the loudspeaker to the error microphone. The signal captured by the reference microphone is denoted as $x(n)$, while the signal captured by the error microphone is denoted as $e(n)$. These signals are processed by the ANC controller to produce a canceling signal $y(n)$, which is emitted by the loudspeaker $f_{LS}$. The loudspeaker output $f_{LS}\{y(n)\}$, after passing through the secondary path $S(z)$, generates the anti-signal denoted by $a(n)$. The relationship is described by: $a(n)=S(z)\ast f_{LS}\{y(n)\}$. Similarly, the reference signal $x(n)$, transmitted through the primary path $P(z)$, produces the primary signal denoted by $d(n)$, which is expressed as: $d(n)=P(z)\ast x(n)$.
The error signal $e(n)$ is the difference between the primary signal $d(n)$ and the anti-signal $a(n)$:
@@
where u and v are the vector representations of the signals $u(n)$ and $v(n)$ such that $\textbf{u}=[u(1),...,u(M)]$ and $\textbf{v}=[v(1),...,v(M)]$. Here, $M$ represents the total number of samples. Typically, $u(n)$ refers to the target signal, while $v(n)$ denotes the estimated signal. A lower NMSE value indicates a better estimation, reflecting a closer alignment between the estimated signal and the target signal. In the context of ANC, typically $u(n)$ is the primary signal $d(n)$, while $v(n)$ will be the anti-signal $a(n)$. A schematic representation of the ANC system is illustrated in Figure 1.
[…]
- NOAS -9.85 -16.53 -18.56 -23.63
+ NOAS -12.77 -17.60 -19.62 -
Table 1: Comparison of NMSE distances for different objectives, with and without NOAS optimization. Measured on DeepASC training set.
Figure 1: Typical feedforward ANC system diagram.
Figure 2: DeepASC Architecture.
Table 2: NMSE focusing on VAD-masked speech-active segments ($\eta^{2}=0.5$).
Method TIMIT ($\downarrow$) WSJ ($\downarrow$) LibriSpeech ($\downarrow$)
DeepANC -9.7 -7.93 -12.63
ARN -8.2 -5.61 -12.29
DeepASC -17.8 -15.56 -17.66
Method Factory ($\downarrow$) Babble ($\downarrow$) WSJ ($\downarrow$)
DeepANC -9.29 -10.94 -8.26
ARN -8.97 -11.17 -10.70
DeepASC -12.09 -13.87 -12.23
Table 4: FLOPs & NMSE comparison for different ANC methods.
Method FLOPs (G) ($\downarrow$) NMSE ($\downarrow$)
DeepANC 7.199 -10.69
ARN 5.281 -11.61
DeepASC 2.419 -13.46
4 Method
We propose a novel architecture that integrates the Mamba framework [49] with a multi-band masking strategy based on Dual-path Mamba blocks [53]. A filter bank splits the input, and each band is processed by an encoder-masker-decoder pipeline. An improved cancellation accuracy is achieved by using a new loss function, which uses a near-optimal anti-signal as ground truth. A diagram of the proposed architecture is shown in Fig. 2.
@@ 4.1 DeepASC Architecture
Let $x(n)$ be the reference signal such that $1\leq n\leq M$. The reference signal $x(n)$ is decomposed into $Q\in\mathbb{N}$ different frequency bands $x_{1}(n),\ldots,x_{Q}(n)$. These frequency bands are evenly divided such that for the maximum frequency $F$, the $i$-th frequency band $x_{i}(n)$ covers the frequency range $\left[(i-1)\frac{F}{Q},i\frac{F}{Q}\right]$ where $1\leq i\leq Q$. In addition to the decomposed bands, the original full-band signal $x(n)$ is included as $x_{0}(n)$. Each band $x_{i}(n)$ (where $0\leq i\leq Q$, the zero index is for the entire unfiltered band) is then processed through its own Masking-Band block (MB-block). Each MB-block comprises an encoder and a masking network that utilize Mamba-based layers. Within each MB-block, the encoder consists of a one-dimensional convolution layer $E_{i}$ with a kernel size $k$ and a stride of $k/2$. The encoder transforms the $i$-th reference signal $x_{i}(n)$ into a two-dimensional latent representation:
@@ 4.2 Optimization Objective
The training protocol for the proposed method consists of two distinct phases: (i) ANC loss minimization, and (ii) near optimal anti-signal fine-tuning optimization. Each phase employs the NMSE loss function (Eq. 2) but with different optimization objectives.
@@
Near Optimal Anti-Signal Optimization (NOAS): A key challenge in formulating ANC as a supervised learning task is designing a training objective function that accounts for the influence of the acoustic paths [6]. In practice, the model output $y(n)$ is nonlinearly transformed by $f_{LS}$ and filtered through $S(z)$, with the goal of minimizing the residual error $e(n)$. However, when $S(z)$ attenuates certain frequency bands that are not attenuated by $P(z)$, conventional loss functions (e.g., Eq. 8) penalize the model despite producing optimal pre-propagation anti-noise. This mismatch introduces misleading gradients, destabilizing training and hindering convergence.
To address this challenge, we propose the NOAS loss function. The NOAS loss symmetrically incorporates the secondary path $S(z)$ on both sides of the NMSE calculation. Specifically, each reference signal $x(n)$ is associated with its NOAS target $y^{*}(n)$. To determine the near-optimal anti-signal $y^{*}(n)$, we employ a gradient descent-based algorithm during a pre-processing stage. This stage operates over each example, solving the following optimization problem for each reference signal $x(n)$ separately:
@@
5 Experiments
5.1 Datasets
The training data is sources from the AudioSet dataset [67], which we encompassed 248 diverse audio classes including hubbub, speech noise, and babble. A total of 22,224 audio samples (approximately 18.5 hours) were standardized to 3 seconds and resampled to 16kHz, following the settings of the ARN method [41]. Of these, 20,000 samples (90%) were used for training and 2,224 for testing. Additional test data were sourced from the NoiseX-92 dataset [68], which includes noise types such as bubble, factory, and engine noise. To evaluate speech generalization, we incorporated test samples from three speech corpora: TIMIT [69] (24 speakers across 8 dialects), LibriSpeech [70] (40 audiobook speakers), and WSJ [71] (8 speakers reading news text).
5.2 Simulator
[…]
5.3 Hyperparameters
An extensive grid search and cross-validation were employed to determine the optimal hyperparameters for each method. The hyperparameter values reported here correspond to the configurations that achieved the best performance in our experimental setup. The DeepASC architecture was trained with $Q=2$, where the full-band employed a medium (M) configuration with 16 layers, and each of the two sub-bands used a small (S) configuration with 8 layers. The bands decomposition filters are generated using the $scipy.signal.firwin$ function and applied to the signal via $torch.conv1d$. The temporal duration $M$ was set to 48,000 samples, corresponding to 3-second audio signals sampled at 16 kHz. The channel dimension $C$ was set to 256, and the kernel size $W$ was defined as 16. A batch size of 2 was used for training the DeepASC architecture. The Adam optimizer [75] was employed with an initial learning rate of $1.5\times 10^{-4}$. A learning rate decay factor of 0.5 was applied every 2 epochs after an initial warm-up period of 30 epochs. Gradient clipping with a threshold of 5 was applied to prevent exploding gradients.
5.4 Baseline Methods
We compared our proposed method against several established ANC techniques, including DeepANC [6], Attentive Recurrent Network (ARN) [41], Filtered-x LMS (FxLMS), and Tangent Hyperbolic Function FxLMS (THF-FxLMS) [19]. All methods were evaluated under identical simulation settings in both linear and nonlinear scenarios, using noise and speech signals. FxLMS, DeepANC, and ARN were implemented and trained by us, employing the same dataset used for our model. DeepANC used 20-ms STFT frames with 10-ms overlap; ARN used 16-ms frames with 8-ms overlap. Our implementations reproduced results consistent with the original papers. These baselines were chosen to ensure a comprehensive comparison across both classical adaptive filtering and recent deep-learning-based ANC paradigms.
6 Results
6.1 Noise Cancellation
Table 5 presents the NMSE results for ANC algorithms under engine, factory, and babble noise using 3-second segments from NoiseX-92. Each model was evaluated with and without nonlinear distortions ($\eta^{2}=\infty$, $0.5$, and $0.1$). For traditional methods (FxLMS and THF-FxLMS), gradient clipping at $1\text{e}^{-4}$ and step sizes of 0.05 (engine), 0.4 (factory), and 0.3 (babble) were used to ensure stability. The results indicate that these methods performed worse than deep learning-based approaches.
Among deep-learning-based models, and without considering the nonlinearity saturation effect, the proposed DeepASC method achieved state-of-the-art performance. Without nonlinearity ($\eta^{2}=\infty$), DeepASC outperformed ARN by 4.29dB, 4.64dB, and 7.26dB for engine, factory, and babble noise, respectivly. With moderate distortion ($\eta^{2}=0.5$), DeepASC yielded respective improvements of 4.36dB, 4.62dB, and 7.13dB. Under severe distortion ($\eta^{2}=0.1$), DeepASC led with a margin of 3.79dB for engine noise, 4.4dB for factory noise, and 5.76dB for babble noise. Figures3(a), 3(b), and 3(c) illustrate that DeepASC consistently outperforms ARN, DeepANC, and FxLMS across all timesteps.
Table 5: Average NMSE ($\downarrow$) in dB for DeepASC and other algorithms across various noise types and nonlinear distortions. Lower values indicate better performance.